28 miles on 1 gallon of gasoline 350 miles on y gallons of gasoline?

what fraction would you use to find 66 2/3% of 72 mentally?
28.5 is what percent of 20?

Answer by pragya.ban@gmail.com(32)

28 miles on 1 gallon of gasoline 350 miles on y gallons of gasoline?

28/1=350/y
Now cross multiply:
28y=350
y=12.5


what fraction would you use to find 66 2/3% of 72?

I would use .666 or .667.


28.5 is what percent of 20?

x%*20=28.5
x%=28.5/20
x=143%

Hope this helps!